OpenStudy (anonymous):

probability: jack is completing a true false test on his laptop that consists of 10 questions / WHAT ARE THE POSSIBLE OUTCOMES !!!!! Answer is :1024

jimthompson5910 (jim_thompson5910):

what's your question? how to get the 1024?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

There are 10 slots and 2 choices per slot, so there are 2*2*2*2*2*2*2*2*2*2 = 2^10 = 1024 different ways to fill it out (ex: one way is TTFTFTTFFT and another is TFFTFTFFTT)